OpenStudy (anonymous):

the angle measures of a triangle are 2x+10deg , 3x deg, and 3x+ 10 deg. find each angle measure.

OpenStudy (anonymous):

The sum (again) is 180. (2x+10) + (3x) + (3x+10) = 180 8x + 20 = 180 8x = 160 x = 20 angle 1 = 2*20 + 10 = 50 deg angle 2 = 3*20 = 60 deg angle 3 = 3*20 + 10 = 70 deg